A vessel w/ a tiny hole in its wall is filled w/ O2 to pressure of 1 atm at 25C. In one minute period, 3.25g of O effuses out through hole into vacuum. The vessel is evacuated and filled with unknown gas at same pressure and temperature. In this case, only 1.96g of the unknown gas effuse out in one minute. Calculate molar mass of unknown gas.
1
Expert's answer
2012-12-03T09:07:24-0500
n of O2 = 3.25/32 = 0.1015mol

the n of unknown gas is the same

n = m/Mw.

Mw = m/n = 1.96/0.1015 = 19